# FRNet: Frustum-Range Networks for Scalable LiDAR Segmentation

### v1

@@ 3.2 Frustum-Range Architecture
3.3 Frustum-Range Operators
Data augmentation enables the model to learn more robust representations. Commonly used techniques in LiDAR segmentation often operate on the global view, including rotation, flipping, scaling, jittering, etc. In this work, we introduce two novel augmentations based on our proposed frustum-range view representation.
FrustumMix. Different from the previous LaserMix [26] and PolarMix [52] augmentations which mix two LiDAR scenes in the point level, we propose a FrustumMix to mix two scenes based on frustum units. Points in the scan will be first grouped into different frustum regions along inclination and azimuth directions. We then randomly split frustum regions equally from two directions and switch the corresponding frustum regions from two scans.
@@ 3.3 Frustum-Range Operators
FPS normal-$\uparrow$ Val normal-$\uparrow$ Test normal-$\uparrow$ Val normal-$\uparrow$ Test normal-$\uparrow$ mIoU normal-$\uparrow$ mAcc normal-$\uparrow$ mIoU normal-$\uparrow$ mAcc normal-$\uparrow$
MinkUNet [10] CVPR’19 Sparse Voxel $\bullet$ $21.7$ M
$\mathbf{9.1}$
$63.7$
$55.0$
$\mathbf{53.1}$
$\mathbf{68.3}$
SPVNAS [47] ECCV’20 Sparse Voxel $\bullet$ $21.8$ M $8.9$
$\underline{66.4}$
$\underline{56.9}$
$48.4$
$61.5$
Cylinder3D [65] CVPR’21 Sparse Voxel $\bullet$ $55.9$ M
$6.2$
$\mathbf{65.9}$
$\mathbf{67.8}$
$\mathbf{76.1}$
$\mathbf{77.9}$
$\mathbf{57.0}$
$\underline{52.9}$
$\underline{64.9}$
RPVNet [54] ICCV’21 Multi-View $\bullet$ $24.8$ M $7.9$
$65.5$
$70.3$
$77.6$
2DPASS [56] ECCV’22 Multi-View $\bullet$ $26.5$ M
$\mathbf{8.4}$
$69.3$
$72.2$
$\mathbf{79.4}$
$80.8$
SphereFormer [27] CVPR’23 Multi-View $\bullet$ $32.3$ M
$4.9$
$67.8$
$74.8$ $78.4$
$\underline{81.9}$
UniSeg [32] ICCV’23 Multi-View $\bullet$ $147.6$ M
$6.9$
$\mathbf{71.3}$
$\mathbf{75.2}$
$\mathbf{83.5}$
RangeNet++ [35] IROS’19 Range View $\bullet$ $50.4$ M
$12.0$
$50.3$
$52.2$
$65.5$
$44.6$
$57.8$
$30.9$
FIDNet [64] IROS’20 Range View $\bullet$ $6.1$ M
$31.8$
$58.9$
$59.5$
$54.1$
$65.4$
$46.4$
CENet [8] ICME’22 Range View $\bullet$ $6.8$ M $33.4$
$62.6$
$64.7$
$55.7$
$66.8$
$50.3$
RangeViT [4] CVPR’23 Range View $\bullet$ $23.7$ M
$10.0$
$60.7$
$64.0$
$75.2$
$53.6$
$66.5$
RangeFormer [24] ICCV’23 Range View $\bullet$ $24.3$ M
$6.2$
$67.6$
$\mathbf{73.3}$
$78.1$
$80.1$
$63.0$
Fast-FRNet Ours Frustum-Range $\bullet$ $7.5$ M
$\mathbf{33.8}$
$67.1$
$\underline{72.5}$
$78.8$ $82.1$
$62.4$
$71.2$ $52.5$ $67.1$
FRNet Ours Frustum-Range $\bullet$ $10.0$ M
$29.1$
$\mathbf{68.7}$
$\mathbf{73.3}$
$\mathbf{79.0}$
$\mathbf{82.5}$
$\mathbf{63.1}$
$\mathbf{72.3}$
$\mathbf{53.5}$
$\mathbf{68.1}$
@@
Sup.-only - FRNet (Ours) $44.9$ $60.4$ $61.8$ $63.1$ $51.9$ $68.1$ $70.9$ $74.6$ $42.4$ $53.5$ $55.1$ $57.0$
@@
Range-Interpolation. Previous range-view methods often suffer from large number of empty pixels, which still exist in FRNet. To alleviate this problem, we propose to interpolate some potential points based on the range image. Specifically, LiDAR points will be first projected to range images as previous works do. We then scan all pixels in the range images and find the empty pixels. For each empty pixel, we will judge whether there are projected points in the pixels along its left and right directions. When both the two pixels have projected points, the interpolated points of this pixel can be calculated by averaging the coordinates of two points. During the training procedure, we also need to generate semantic labels for the interpolated points, which will face two cases. If both right and left points have the same semantic labels, we will set the semantic label of the interpolated point to be the same as them. For another case with different semantic labels, we consider the interpolated points as boundaries of the object, and it will be set to an ignored label, which will not be used to calculate loss.
4 Experiments
4.1 Settings
@@
Datasets. We conduct experiments on four popular LiDAR segmentation datasets. SemanticKITTI [5] consists of $22$ sequences with a total of $43,552$ LiDAR scans. Officially, sequences $00$ to $07$ and $09$ to $10$ ($19,130$ scans) are used for training, sequence $08$ ($4,071$ scans) for validation, and sequences $11$ to $21$ ($20,351$ scans) for online testing with $19$ semantic classes. nuScenes [13] provides $1,000$ driving scenes captured with much sparser points by 32-beam LiDAR sensors, and 16 classes are evaluated during inference. SemanticPOSS [37] is a more challenging benchmark collected by Peking University with much smaller and sparser point clouds. It consists of $2,988$ LiDAR scenes with a large number of dynamic instances. Officially, it is divided into $6$ parts, and part $2$ is used to evaluate the performance with $13$ categories. ScribbleKITTI [49] shares the same scenes with SemanticKITTI [5] but only provides weak annotations with line scribbles, which means around $8.06\%$ labeled points available during training.
@@
Evaluation Metrics. In line with standard protocols, we utilize Intersection-over-Union (IoU) and Accuracy (Acc) for each class $i$ and compute the mean Intersection-over-Union (mIoU) and mean Accuracy (mAcc) to access the performance. IoU and Acc can be calculated as:
$\text{IoU}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}+\text{FN}_{i}}~{},~{}~{}\text{Acc}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}}~{},$ (9)
@@
where $\text{TP}_{i},\text{TP}_{i},\text{FN}_{i}$ are true-positives, false-positives, and false-negatives for class $i$, respectively. For robustness evaluation, we follow the practice of Robo3D [25] to use Corruption Error (CE) and Resilience Rate (RR) to evaluate the robustness of FRNet. The CE and RR can be calculated as:
$\text{CE}_{i}=\frac{\sum_{l=1}^{3}(1-\text{IoU}_{i,l})}{\sum_{l=1}^{3}(1-\text{IoU}_{i,l}^{\text{base}})}~{},~{}~{}\text{RR}_{i}=\frac{\sum_{l=1}^{3}\text{IoU}_{i,l}}{3\times\text{IoU}_{\text{clean}}}~{},$ (10)
@@
where $\text{IoU}_{i,l}$ is the task-specific IoU; $\text{IoU}_{i,l}^{\text{base}}$ and $\text{IoU}_{\text{clean}}$ are scores of the baseline model and scores on the clean evaluation set.
@@
Implementation Details. We implement FRNet using the MMDetection3D [11] framework and utilize CENet [8] as our backbone to extract frustum features. We set the shape of the 2D representation to $64\times 512$ for SemanticKITTI [5] and ScribbelKITTI [49] and $32\times 480$ for nuScenes [13] and SemanticPOSS [37], respectively. We choose AdamW [34] as our default optimizer and utilize the OneCycle scheduler [45] with $lr=0.01$. The models are trained for $50$ epochs on SemanticKITTI [5] and ScribbelKITTI [49], and $80$ epochs on nuScenes [13] and SemanticPOSS [37], respectively, with a batch size of $2$ with $4$ GPUs.
4.2 Comparative Study
@@
Comparisons to State of the Arts. Firstly, we compare the proposed FRNet with state-of-the-art LiDAR segmentation methods from different representations on both fully-supervised and weakly-supervised benchmarks. Tab. 1 summarizes all the results and the FPS of all the models is evaluated with a single GeForce RTX 2080Ti GPU for fair comparison. Fast-FRNet takes resnet18 [19] as the 2D backbone and the shapes of 2D representation on four datasets are $32\times 360$. Compared with multi-view methods such as SphereFormer [27] and UniSeg [32], although FRNet does not have a significant advantage in performance, its parameters and FPS are far superior to theirs. Compared with voxel-view and range-view methods, FRNet achieves appealing performance while still maintaining high efficiency, and Fast-FRNet achieves the fastest inference speed with only a little drop in performance. Fig. 6 also shows qualitative results among state-of-the-art methods, including CENet [8] and RangeViT [4] on the val set of SemanticKITTI [5]. FRNet reduces prediction errors over large areas, especially for vegetation and parking.
Label-Efficient LiDAR Segmentation. Recently, semi-supervised learning has been explored in LiDAR segmentation. In this work, we follow the paradigm of LaserMix [26] to conduct FRNet on the semi-supervised segmentation task. Here, we use FRNet as the backbone and the proposed FrustumMix to mix labeled and unlabeled. Tab. 2 summarizes the results of recent methods. Results show that our FRNet backbone achieves appealing performance among state-of-the-art methods. Meanwhile, we compare the proposed FrustumMix with LaserMix among different backbones, including Cylinder3D [65] and FRNet. Results show that the proposed FrustumMix outperforms LaserMix, even in voxel representation.
Robustness. Robustness reflects the model’s ability to generalization to a certain extent, which is particularly important in autonomous driving, especially in extreme weather conditions. To this end, we also evaluate the robustness of FRNet using the noisy datasets SemanticKITTI-C and nuScenes-C introduced in Robo3D [25], which include 8 different conditions and each condition has 3 different levels. The models trained with SemanticKITTI [5] and nuScenes [13] are used to evaluate the performance on SemanticKITTI-C and nuScenes-C [25], respectively. As shown in Tab. 3, FRNet demonstrates strong robustness, especially under extreme weather conditions.
@@
Method Venue SemKITTI-C nuScenes-C
mCE normal-$\downarrow$ mRR normal-$\uparrow$ mCE normal-$\downarrow$ mRR normal-$\uparrow$
MinkUNet [10] CVPR’19 $100.0$
$\mathbf{81.9}$
$100.0$
$74.4$
Cylinder3D [65] CVPR’21
$103.3$
$80.1$
$111.8$
$72.9$
2DPASS [56] ECCV’22
$106.1$
$77.5$
$\mathbf{98.6}$
$75.2$
FIDNet [64] IROS’20
$113.8$
$77.0$
$122.4$
$73.3$
CENet [8] ICME’22
$103.4$
$81.3$
$112.8$
$76.0$
FRNet Ours
$\mathbf{96.8}$
$80.0$
$\mathbf{98.6}$
$\mathbf{77.5}$
Figure 5: Ablation study on different configurations in FRNet. Tables show the comparisons of: a) Different range-view backbones on the val set of SemanticKITTI [5]. b) Different mixing strategies on the val set of nuScenes [13]. c) Different mixing strategies on the val set of SemanticKITTI [5]. Color-filled boxes denote mAcc while empty boxes denote mIoU in (b) and (c).
@@
Figure 6: Qualitative results among state-of-the-art LiDAR segmentation methods [8, 4]. To highlight the differences compared with groundtruth, the correct and incorrect predictions are painted in gray and violet, respectively. Each point cloud scene is visualized under both the bird’s-eye-view and range-view.
4.3 Ablation Study
In this section, we will discuss each component in FRNet on the val set of SemanticKITTI [5] and nuScenes [13].
Model Analysis. We first analyze different components in FRNet, including the Frustum Feature Encoder, Frustum-Point (FP) Fusion, Fusion Head, Frustum-level Supervision, and Range-Interpolation. The results are summarized in Tab. 4. As the baseline of FRNet, the Frustum Feature Encoder module extracts per-point features in the frustum region and predicts semantic labels of each point with the supervision of point-level labels, and the performance is $62.7\%$ and $72.8\%$ on SemanticKITTI [5] and nuScenes [13], respectively. When we use the FP Fusion module to update per-point features hierarchically, the performance improves by $1.6\%$ and $0.8\%$ on the two datasets, respectively. With the frustum-level supervision, the performance further improves by $1.9\%$ and $2.8\%$. Next, we utilize a Fusion Head module to fuse different level features for predicting, and it improves by $0.8\%$ on both datasets. To alleviate the empty pixels in 2D representations, we propose a Range-Interpolation method to generate some potential points. Results show that this augmentation can also help to improve the performance by $0.6\%$ and $0.5\%$. Finally, following most works, we also use Test Time Augmentation (TTA) during inference, and it will bring an improvement of $1.1\%$ and $1.3\%$, respectively.

### v2

@@ III-B Frustum-Range Architecture
Fusion Head Module. FRNet aims to assign a category for each point in the point cloud, which requires point-level features for end-to-end label prediction. Although output point features at the last layer of the frustum-point fusion module can yield promising performance, we observe this is not the optimal solution. Features from different levels have different context-aware information. Lower-level features are geometric while higher-level features are semantically rich. Leveraging these different level features can enhance performance. Drawing inspiration from FIDNet [55] and CENet [58], we first combine all point and frustum features from each stage of the backbone and fuse them to reduce channel dimension:
$\mathcal{F}_{p}^{\text{out}}=\text{MLP}([\mathcal{F}_{p}^{1};...;\mathcal{F}_{p}^{V}])~{},$ (7)
$\mathcal{F}_{f}^{\text{out}}=\text{Conv}([\mathcal{F}_{f}^{1};...;\mathcal{F}_{f}^{V}])~{},$ (8)
where $V$ is the number of stages in the backbone. Fusing backbone features with different receptive fields helps obtain more context-aware and semantic information. Additionally, point features $\mathcal{F}_{p}^{\text{out}}$ contains local information from neighborhood frustum regions, while frustum features $\mathcal{F}_{f}^{\text{out}}$ contains global information within the corresponding frustum region. Meanwhile, $\mathcal{F}_{p}^{0}$ from the frustum feature encoder module constructs graph structures within the frustum region and between neighborhood frustum regions. Progressively fusing these features helps extract features from local to global, and from geometric to semantic, resulting in more accurate predictions. To fuse $\mathcal{F}_{f}^{\text{out}}$ and $\mathcal{F}_{p}^{\text{out}}$, we first back-project frustum features to point-level and use an MLP layer ensure the same channels as $\mathcal{F}_{p}^{\text{out}}$. Then, an addition operation is applied to fuse the two features. The same operation is utilized to fuse the new features with $\mathcal{F}_{p}^{0}$. The overall operation can be formulated as:
$\mathcal{F}_{\text{logit}}=\text{MLP}(\text{MLP}(\text{Inflat}(\mathcal{F}_{f}^{\text{out}}))+\mathcal{F}_{p}^{\text{out}})+\mathcal{F}_{p}^{0}~{}.$ (9)
Here, $\mathcal{F}_{\text{logit}}$ is used to generate the final semantic scores with a linear head for the point over the entire point cloud.
Optimization. In the conventional setup, a point-level loss function is commonly employed to supervise the final semantic scores. However, in this work, we propose optimizing frustum features with additional auxiliary heads, utilizing the frustum-level loss function. This approach requires generating pseudo frustum labels. Typically, a frustum region encompasses multiple points, which may not all belong to the same category. To address this, we adopt a statistical approach inspired by sparse-voxel methods [16]. We generate a pseudo label for each frustum by first counting the frequency of each category existing in the given frustum region, and then choosing the category with the highest frequency as the pseudo label of the frustum. Finally, the frustum-level features are supervised using the generated pseudo labels. The overall loss function can be formulated as follows:
$\mathcal{L}=\mathcal{L}_{p}+\lambda\mathcal{L}_{f}~{},$ (10)
where $\mathcal{L}_{p}$ and $\mathcal{L}_{f}$ represent the point-level and frustum-level loss functions, respectively, and $\lambda$ is a hyperparameter that controls the weight of frustum supervision. $\mathcal{L}_{p}$ is calculated by cross-entropy loss, while $\mathcal{L}_{f}$ consists of cross-entropy loss, Lovasz-Softmax loss [61], and boundary loss [62].
III-C Frustum-Range Operators
Data augmentation enables the model to learn more robust representations. Common augmentations in LiDAR segmentation typically operate on the global view, encompassing rotation, flipping, scaling, jittering, etc. In this work, we introduce two novel augmentations based on our proposed frustum-range view representation.
FrustumMix. Previous mixing strategies [44, 45] involve swapping two LiDAR scans at the point level, which may disrupt the internal structure within the frustum region. In this work, we propose FrustumMix, a method to swap frustum regions from two scenes. Specifically, given two LiDAR point clouds $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, we randomly split the point clouds into $M$ non-overlapping frustum regions along the inclination or azimuth direction via Eq. 1, denoted as $\mathcal{P}_{1}=\{\mathcal{A}_{1}^{1},...,\mathcal{A}_{1}^{M}\}$ and $\mathcal{P}_{2}=\{\mathcal{A}_{2}^{1},...,\mathcal{A}_{2}^{M}\}$. We then generate a new LiDAR scan $\hat{\mathcal{P}}$ by alternately swapping frustum regions from $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, as follows: $\hat{\widetilde{P}}=\mathcal{A}_{1}^{1}\cup\mathcal{A}_{2}^{2}\cup\mathcal{A}_{1}^{3}\cup\mathcal{A}_{2}^{4}\cup\cdots$. FrustumMix enhances context-awareness between frustums while preserving the invariant structure within the frustum.
@@ III-C Frustum-Range Operators
FPS $\uparrow$ Val $\uparrow$ Test $\uparrow$ Val $\uparrow$ Test $\uparrow$ mIoU $\uparrow$ mAcc $\uparrow$ mIoU $\uparrow$ mAcc $\uparrow$
MinkUNet [17] CVPR’19 Sparse Voxel $\bullet$ $21.7$ M $9.1$ $62.8$ $63.7$ $75.8$ $55.0$ $\mathbf{53.1}$ $\mathbf{68.3}$
SPVNAS [15] ECCV’20 Sparse Voxel $\bullet$ $21.8$ M $8.9$ $62.5$ $66.4$ $74.4$ $56.9$ $48.4$ $61.5$
Cylinder3D [16] CVPR’21 Sparse Voxel $\bullet$ $55.9$ M $6.2$ $65.9$ $67.8$ $\mathbf{76.1}$ $\mathbf{77.9}$ $\mathbf{57.0}$ $52.9$ $64.9$
PVKD [65] CVPR’22 Sparse Voxel $\bullet$ $14.1$ M $\mathbf{13.2}$ $\mathbf{66.4}$ $\mathbf{71.2}$ $76.0$
KPConv [7] ICCV’19 Raw Points $\bullet$ $18.3$ M $3.2$ $61.3$ $58.8$
RandLA-Net [8] CVPR’20 Raw Points $\bullet$ $1.2$ M $1.9$ $57.1$ $53.9$
PTv2 [66] NeurIPS’22 Raw Points $\bullet$ $12.8$ M $4.7$ $\mathbf{70.3}$ $\mathbf{72.6}$ $\mathbf{80.2}$ $\mathbf{82.6}$
WaffleIron [11] ICCV’23 Raw Points $\bullet$ $6.8$ M $\mathbf{7.4}$ $68.0$ $70.8$ $79.1$
RPVNet [21] ICCV’21 Multi-View $\bullet$ $24.8$ M $7.9$ $65.5$ $70.3$ $77.6$
2DPASS [67] ECCV’22 Multi-View $\bullet$ $26.5$ M $\mathbf{8.4}$ $69.3$ $72.2$ $\mathbf{79.4}$ $80.8$
SphereFormer [68] CVPR’23 Multi-View $\bullet$ $32.3$ M $4.9$ $67.8$ $74.8$ $78.4$ $81.9$
UniSeg [23] ICCV’23 Multi-View $\bullet$ $147.6$ M $6.9$ $\mathbf{71.3}$ $\mathbf{75.2}$ $\mathbf{83.5}$
RangeNet++ [30] IROS’19 Range View $\bullet$ $50.4$ M $12.0$ $50.3$ $52.2$ $65.5$ $44.6$ $57.8$ $30.9$
FIDNet [55] IROS’20 Range View $\bullet$ $6.1$ M $31.8$ $58.9$ $59.5$ $71.4$ $54.1$ $65.4$ $46.4$
CENet [58] ICME’22 Range View $\bullet$ $6.8$ M $33.4$ $62.6$ $64.7$ $73.3$ $55.7$ $66.8$ $50.3$
RangeViT [31] CVPR’23 Range View $\bullet$ $23.7$ M $10.0$ $60.7$ $64.0$ $75.2$ $53.6$ $66.5$
RangeFormer [32] ICCV’23 Range View $\bullet$ $24.3$ M $6.2$ $67.6$ $\mathbf{73.3}$ $78.1$ $80.1$ $63.0$
Fast-FRNet Ours Frustum-Range $\bullet$ $7.5$ M $\mathbf{33.8}$ $67.1$ $72.5$ $78.8$ $82.1$ $62.4$ $71.2$ $52.5$ $67.1$
FRNet Ours Frustum-Range $\bullet$ $10.0$ M $29.1$ $\mathbf{68.7}$ $\mathbf{73.3}$ $\mathbf{79.0}$ $\mathbf{82.5}$ $\mathbf{63.1}$ $\mathbf{72.3}$ $\mathbf{53.5}$ $\mathbf{68.1}$
RangeInterpolation. Previous range-view methods [58, 55, 32] often encounter a large number of empty pixels due to the sparsity of LiDAR points, resulting in noisy information. To mitigate this problem, we introduce RangeInterpolation, a technique that utilizes the range image to reconstruct surfaces in the LiDAR scan. Specifically, the point cloud $\mathcal{P}$ is first projected onto the range image $\mathcal{R}\in\mathbb{R}^{H\times W\times(3+L)}$. For empty pixel position $(u,v)$ in the range image $\mathcal{R}$, we establish a window with size $m\times n$ centered at $(u,v)$, where $m$ and $n$ are odd. Our aim is to interpolate the point falling into $(u,v)$ using the surrounding range information within the predefined window. Considering that not all pixels in the window contain valid values, we aggregate the information from non-empty pixels within the window for point interpolation using an average operation. During the training procedure, the semantic label of the interpolated point is determined via threshold-guided voting within the window. Specifically, we identify the most frequent category within the window and its frequency. If the frequency is below a threshold, we classify the interpolated point as located at the boundary and assign the label as ignored, which will not contribute to loss calculation. Otherwise, the label is determined by the category with the highest frequency.
IV Experiments
In this section, we demonstrate the scalability and robustness of the proposed method. We first introduce the datasets, metrics, and detailed implementations we used. Subsequently, we show that FRNet achieves an optimal trade-off between efficiency and accuracy when compared to state-of-the-art methods across various task setups. Finally, we conduct a series of ablation studies to analyze each component in FRNet.
@@ IV Experiments
Sup.-only - FRNet $44.9$ $60.4$ $61.8$ $63.1$ $51.9$ $68.1$ $70.9$ $74.6$ $42.4$ $53.5$ $55.1$ $57.0$
PolarMix [45] NeurIPS’22 $50.1$ $60.9$ $62.0$ $63.8$ $55.6$ $69.6$ $71.0$ $73.8$ $43.2$ $55.0$ $56.1$ $57.3$
@@
Figure 5: Class-wise LiDAR segmentation results of FRNet and the baseline model on the test set of SemanticKITTI [28].
IV-A Experimental Settings
@@
Datasets. We conduct a comprehensive experiment on four popular LiDAR segmentation datasets. SemanticKITTI [28] is a large-scale outdoor dataset that comprises $22$ sequences collected from various scenes in Karlsruhe, Germany, using a Velodyne HDL-64E LiDAR sensor. Officially, sequences $00$ to $07$ and $09$ to $10$ ($19,130$ scans) are used for training, sequence $08$ ($4,071$ scans) for validation, and sequences $11$ to $21$ ($20,351$ scans) for online testing. The original dataset is annotated with $28$ classes, though only $19$ merged classes are used for single-scan evaluation. The vertical FOV is $[-25^{\circ},3^{\circ}]$. nuScenes [43] is a multimodal dataset that is widely used in autonomous driving scenarios. It contains LiDAR point clouds captured around streets in Boston and Singapore, provided by 32-beam LiDAR sensors. The dataset includes $1,000$ driving scenes with sparser points. It is annotated with $32$ classes and only $16$ semantic categories are used for evaluation. The vertical FOV is $[-30^{\circ},10^{\circ}]$. SemanticPOSS [64] is a more challenging benchmark collected at Peking University and features much smaller and sparser point clouds. It consists of $2,988$ LiDAR scenes with numerous dynamic instances. Officially, it is divided into $6$ sequences, with sequence $2$ used for evaluation with $13$ merged categories. The vertical FOV is $[-16^{\circ},7^{\circ}]$. ScribbleKITTI [63] shares the same scenes with SemanticKITTI [28] but only provides weak annotations with line scribbles. It contains $189$ million labeled points, with approximately $8.06\%$ labeled points available during training. Additionally, we conduct a comprehensive out-of-distribution robustness evaluation experiment on SemanticKITTI-C and nuScenes-C from the Robo3D [71] benchmark. Each dataset contains eight corruption types, including “fog”, “wet ground”, “snow”, “motion blur”, “beam missing”, “crosstalk”, “incomplete echo”, and “cross sensor”.
@@
Evaluation Metrics. In line with standard protocols, we utilize Intersection-over-Union (IoU) and Accuracy (Acc) for each class $i$ and compute the mean Intersection-over-Union (mIoU) and mean Accuracy (mAcc) to assess performance. IoU and Acc can be calculated as follows:
$\text{IoU}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}+\text{FN}_{i}}~{},\text{Acc}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}}~{},$ (11)
@@
where $\text{TP}_{i},\text{FP}_{i},\text{FN}_{i}$ are true-positives, false-positives, and false-negatives for class $i$, respectively. For robustness evaluation, we follow the practice of Robo3D [71] to use Corruption Error (CE) and Resilience Rate (RR) to evaluate the robustness of FRNet. The CE and RR are calculated as follows:
$\text{CE}_{i}=\frac{\sum_{l=1}^{3}(1-\text{IoU}_{i,l})}{\sum_{l=1}^{3}(1-\text{IoU}_{i,l}^{\text{base}})}~{},\text{RR}_{i}=\frac{\sum_{l=1}^{3}\text{IoU}_{i,l}}{3\times\text{IoU}_{\text{clean}}}~{},$ (12)
@@
where $\text{IoU}_{i,l}$ represents the task-specific IoU. $\text{IoU}_{i,l}^{\text{base}}$ and $\text{IoU}_{\text{clean}}$ denote scores of the baseline model and scores on the “clean” evaluation set, respectively.
@@
Implementation Details. We implement FRNet using the popular MMDetection3D [72] codebase and employ CENet [58] as the 2D backbone to extract frustum features. The 2D representation shape is set to $64\times 512$ for SemanticKITTI [28] and ScribbelKITTI [63], and $32\times 480$ for nuScenes [43] and SemanticPOSS [64], respectively. For optimization, We choose AdamW [73] as our default optimizer with an initial learning rate of $0.01$. We utilized the OneCycle scheduler [74] to dynamically adjust the learning rate during training. All models are trained using four GPUs for $50$ epochs on SemanticKITTI [28] and ScribbelKITTI [63], and $80$ epochs on nuScenes [43] and SemanticPOSS [64], respectively. The batch size is set to $4$ for each GPU. Furthermore, we implement a faster variant, dubbed Fast-FRNet, which features a smaller 2D representation resolution and scaled-down 2D backbone. Specifically, the size of the 2D representation is adjusted to $32\times 360$ for all datasets, and the 2D backbone is a variant of ResNet18 [75], following the design of CENet [58].
IV-B Comparative Study
@@
Comparisons to State of the Arts. We begin by comparing FRNet with state-of-the-art LiDAR segmentation methods across different representations, including sparse-voxel, point-view, multi-view, and range-view representations, on both fully-supervised and weakly-supervised benchmarks. Tab. I provides a summary of all the results. We evaluate the Frame Per Second (FPS) on a single GeForce RTX 2080Ti GPU for fair comparison. For SemanticKITTI [28] and nuScenes [43], we report mIoU scores on both the val and test sets. For ScribbleKITTI [63] and SemanticPOSS [64], which share the same data on val and test sets, we report both mIoU and mAcc scores on the val set. Compared with multi-view methods, although FRNet does not exhibit a significant advantage in performance, its parameters and FPS are notably superior. In comparison with sparse-voxel-view and range-view methods, FRNet achieves appealing performance while still maintaining high efficiency. Additionally, Fast-FRNet demonstrates the fastest inference speed with only a minor drop in performance.
@@
# Method Venue mCE $\downarrow$ mRR $\uparrow$ mIoU $\uparrow$ Fog Weg Snow Motion Beam Cross Echo Sensor
SemanticKITTI-C MinkUNet [17] CVPR’19 $100.0$ $81.9$ $62.8$ $\mathbf{55.9}$ $54.0$ $53.3$ $32.9$ $56.3$ $\mathbf{58.3}$ $54.4$ $46.1$
SPVCNN [15] ECCV’20 $100.3$ $\mathbf{82.2}$ $62.5$ $55.3$ $54.0$ $51.4$ $34.5$ $56.7$ $58.1$ $54.6$ $46.0$
FIDNet [55] IROS’20 $113.8$ $77.0$ $58.8$ $43.7$ $51.6$ $49.7$ $40.4$ $49.3$ $49.5$ $48.2$ $29.9$
Cylinder3D [16] CVPR’21 $103.3$ $80.1$ $63.4$ $37.1$ $57.5$ $46.9$ $52.5$ $57.6$ $56.0$ $52.5$ $46.2$
2DPASS [67] ECCV’22 $106.1$ $77.5$ $64.6$ $40.5$ $60.7$ $48.5$ $\mathbf{57.8}$ $58.8$ $28.5$ $55.8$ $50.0$
CENet [58] ICME’22 $103.4$ $81.3$ $62.6$ $42.7$ $57.3$ $53.6$ $52.7$ $55.8$ $45.4$ $53.4$ $45.8$
WaffleIron [11] ICCV’23 $109.5$ $72.2$ $66.0$ $45.5$ $58.6$ $49.3$ $33.0$ $59.3$ $22.5$ $\mathbf{58.6}$ $\mathbf{54.6}$
FRNet Ours $\mathbf{96.8}$ $80.0$ $\mathbf{67.6}$ $47.6$ $\mathbf{62.2}$ $\mathbf{57.1}$ $56.8$ $\mathbf{62.5}$ $40.9$ $58.1$ $47.3$
nuScenes-C MinkUNet [17] CVPR’19 $100.0$ $74.4$ $75.8$ $53.6$ $73.9$ $40.4$ $\mathbf{73.4}$ $\mathbf{68.5}$ $26.6$ $\mathbf{63.8}$ $\mathbf{51.0}$
SPVCNN [15] ECCV’20 $106.7$ $74.7$ $74.4$ $59.0$ $72.5$ $41.1$ $58.4$ $65.4$ $36.8$ $62.3$ $49.2$
FIDNet [55] IROS’20 $122.4$ $73.3$ $71.4$ $64.8$ $68.0$ $59.0$ $48.9$ $48.1$ $57.5$ $48.8$ $23.7$
Cylinder3D [16] CVPR’21 $111.8$ $72.9$ $76.2$ $59.9$ $72.7$ $58.1$ $42.1$ $64.5$ $44.4$ $60.5$ $42.2$
2DPASS [67] ECCV’22 $\mathbf{98.6}$ $75.2$ $\mathbf{77.9}$ $64.5$ $\mathbf{76.8}$ $54.5$ $62.0$ $67.8$ $34.4$ $63.2$ $45.8$
CENet [58] ICME’22 $112.8$ $76.0$ $73.3$ $67.0$ $69.9$ $61.6$ $58.3$ $50.0$ $\mathbf{60.9}$ $53.3$ $24.8$
WaffleIron [11] ICCV’23 $106.7$ $72.8$ $76.1$ $56.1$ $73.9$ $49.6$ $59.5$ $65.2$ $33.1$ $61.5$ $44.0$
FRNet Ours $\mathbf{98.6}$ $\mathbf{77.5}$ $77.7$ $\mathbf{69.1}$ $76.6$ $\mathbf{69.5}$ $54.5$ $68.3$ $41.4$ $58.7$ $43.1$
Label-Efficient LiDAR Semantic Segmentation. Recently, semi-supervised learning has been explored in LiDAR segmentation. In this work, we adopt the approach of LaserMix [44] to apply FRNet to the semi-supervised segmentation task, as shown in Tab. II. Firstly, we utilize Cylinder3D [16] as the backbone and leverage the proposed FrustumMix technique to mix labeled and unlabeled data for consistency learning. Notably, FrustumMix demonstrates superior performance compared to the LaserMix [44] method. Subsequently, we employ FRNet as the backbone and assess various strategies for mixing labeled and unlabeled scans, including PolarMix [45], LaserMix [44], and our FrustumMix. The Sup.-only baseline results are obtained without FrustumMix and RangeInterpolation as data augmentation techniques. Results show that FrustumMix consistently achieves the best scores across all data splits, particularly with limited annotations.
Out-of-Distribution Robustness. Robustness, which reflects a model’s ability to generalize under different conditions, is crucial for automotive security systems, particularly in extreme weather conditions [76, 77]. To evaluate the out-of-distribution robustness of FRNet, we utilize the noisy datasets SemanticKITTI-C and nuScenes-C introduced in Robo3D [71]. The models trained on SemanticKITTI [28] and nuScenes [43] are directly applied to evaluate the performance on SemanticKITTI-C and nuScenes-C [71], respectively. As shown in Tab. III, FRNet achieves promising performance across most corruption types and demonstrates superiority over recent works employing various LiDAR representations, including sparse voxel [17, 15, 16], range view [55, 58], multi-view fusion [67], and raw points [11].
@@
Figure 6: Qualitative results among state-of-the-art LiDAR segmentation methods [55, 58, 31] on the val set of SemanticKITTI [28]. To highlight the differences compared with groundtruth, the correct and incorrect predictions are painted in gray and red, respectively. Best viewed in colors and zoomed-in for details.